How do I calculate this in excel?

Calculate the modified B/C ratio for a government project that is projected to have the following cash flows: costs of $4,400,000 per year; benefits of $5,200,000 per year; dis-benefits of $420,000 per year.
Answer #1

B/C ratio = expected benefit from the project / cost of project

= 4,780,000 / 4,400,000

= 1.08

expected benefit from the project = benefits - dis-benefit

= 5,200,000 - 420,000

= 4,780,000